OpenStudy (vheah):

You are given four springs : 2.5 N/m, 5 N/m, 7.5N/m, and 10 N/m How do you solve for the smallest equivalent stiffness that can be made using only three of these springs?

OpenStudy (aliqanber):

For smallest equivalent stiffness, choose the 3 springs with the lowest stiffness. So, choose: 2.5,5,7.5 N/m. Now, you can either orient these springs in parallel or you can put them in series. If in parallel: \[k _{eqv} = k _{1}+k _{2}+k _{3} = 2.5+5+7.5 = 15 N/m\] If in series: \[\frac{ 1 }{ k _{eqv} }=\frac{ 1 }{ k _{1} }+\frac{ 1 }{ k _{2} }+\frac{ 1 }{ k _{3} }=\frac{ 1 }{ 2.5 }+\frac{ 1 }{ 5 }+\frac{ 1 }{ 7.5 }=0.73333\] \[k _{eqv}=\frac{ 1 }{ 0.73333 } = 1.3636N/m\] So, the springs with stiffnesses 2.5,5,7.5 N/m in a series configuration, will give the lowest possible stiffness of about 1.36 N/m.